OpenStudy (anonymous):

Integrate: (sinx+2cosx)^2

OpenStudy (johnweldon1993):

\[\large \int (sin(x) + 2cos(x))^2 = \int (sin^2(x) + 4sin(x)cos(x) + 4cos^2(x)\] \[\large \int sin^2(x)dx + \int 4sin(x)cos(x)dx + \int 4cos^2(x)dx\]

OpenStudy (johnweldon1993):

\[\large \int sin^2(x)dx + 4\int sin(x)cos(x)dx + 4\int cos^2(x)dx\] Doe that make it look any better?

OpenStudy (anonymous):

thank you sorry my phone wouldn't let me see your reply yesterday! thank you very much